OpenStudy (anonymous):

Find the greatest common factor of 270 and 360

OpenStudy (anonymous):

Divide 360 by 270 and get the remainder first...

OpenStudy (anonymous):

ok, 1.3

OpenStudy (anonymous):

Or second thing you can do is to take prime factors of 270 and 360..

OpenStudy (anonymous):

they are both fairly obviously divisible by 10 it is also clear that they are both divisible by 9? 27 is, as is 36 so we know they are both divisible by 90

OpenStudy (anonymous):

i have to break it down.

OpenStudy (anonymous):

into exponents.

OpenStudy (anonymous):

therefore the greatest common factor is at least 90 \[360\div 90=4,270/90=3\] so 90 is the answer

OpenStudy (cwrw238):

i did this one a while ago find prime factors and look for common values 270 = 2*3*3*3*5 360 = 2*2*2*3*3*5 GCF = 2*3*3*5 = 90

OpenStudy (anonymous):

"(Give the answer in the numerical form in the top box and in exponential form by filling in the boxes for exponents.)"

OpenStudy (anonymous):

No. do not divide it properly see do like this:  Get the remainder 90?? This is not the answer yet.. Now divide the divisor 270 by the remainder ie 90... If it is wholly divisible then 90 is the GCF...